OpenStudy (anonymous):

Which of the following are point-slope equations of the line going through (3, 6) and (1, 2)? All that apply. A. y - 6 = 2(x - 3) B. y - 2 = 2(x - 1) C. y - 2 = 1/2 (x - 1) D. y - 6 = 1/2 (x + 3) E. y - 6 = 1/2 (x - 3) F. y - 2 = 2(x + 1)

OpenStudy (anonymous):

Tai!!! :D

OpenStudy (paxpolaris):

if you get the slope 'm' right: you can write the equation as: \[\huge y-k=m \left( x-h \right)\]for any point (h,k) on the line

OpenStudy (paxpolaris):

\[m={6-2 \over 3-1}=2\]

OpenStudy (paxpolaris):

A,B

OpenStudy (anonymous):

Thanks Pax . I got the A , but the B i was still working on. :)

OpenStudy (paxpolaris):

if m is 2: y-k =2(y-h) for A. (h,k)= (3,6) plugging in (3. 6 ) gives you 0=0 for B. (h,k)= (1,2) plugging in (1,2 ) gives you 0=0